# Unified Software Engineering agent as AI Software Engineer

### v1

@@ Abstract.
*Joint first authors, ordered alphabetically.
[…]
ccs: Software and its engineering Automatic programmingccs: Software and its engineering Software maintenance toolsccs: Software and its engineering Software evolution
@@ 1. Introduction
Large language models (LLMs) have shown promise in coding, reasoning, and problem solving. The model capability itself has improved significantly in terms of reasoning tasks, by training the models to always use reasoning paradigms such as chain of thoughts, without needing further prompting, as shown by GPT o1 (OpenAI, 2024b) and its descendant models such as o3. On top of improved foundation models, the software engineering community is distilling known best practices for downstream tasks into amplifiers for LLMs through agentic systems (Yao et al., 2023). An LLM agent for software engineering invokes various interfaces including testing and analysis tools autonomously driven by an LLM reasoning agent.
@@
While agentic systems inspire researchers (Yao et al., 2023; Antoniades et al., 2024), start-ups (Orwall and contributors, 2025; AIDER Team, 2024) and major companies (Wadhwa et al., 2024; Rondon et al., 2025), we currently observe a strong specialization in the emerging technologies among the LLM agents for software. Agentless (Xia et al., 2024) performs program repair, Large Language Monkeys (Brown et al., 2024) produce unit tests, ExecutionAgent (Bouzenia and Pradel, 2024) performs a project setup, and so forth. These specialized agents are reasonable efforts, especially as they require specialized metrics and datasets. However, with many agents comes the need to manage and maintain them as part of the future development environments!
@@
We propose a unified Software Engineering agent (USEagent) representing a consolidated agentic capability for software engineering tasks. Individual SE tasks could be coordinated to result in a more effective ensemble of tasks: better test-generation should benefit reproduction in program repair, review of patches should lead to better generated code, etc. With an integrated software engineering agent, it becomes much more feasible to co-opt it in future development environments (A Roychoudhury and Ray, 2025).
@@
As a foundation, we need a unified dataset of challenging SE tasks to test our unified agentic capabilities, for which we build Unified Software Engineering Bench, or USEbench for short. Recently, the SWE-bench dataset (Jimenez et al., 2024) has been proposed that captures a set of GitHub issues depicted in natural language and requires bug fixes or feature additions in software projects. Thus the SWE-bench dataset is replete with challenges in software maintenance for various real-life software projects. Our proposed benchmark USEbench is a meta-benchmark composing multiple software engineering tasks (such as code generation, program repair, test generation) behind a unified application programming interface (API). Building a unified API is key, since it allows us to start thinking of a unified agentic capability which can handle different software engineering tasks. After constructing USEbench, we expand two popular agentic systems, one from industry and one from academia, AutoCodeRover (Zhang et al., 2024b) and OpenHands CodeActAgent (Wang et al., 2024b, a) to solve the tasks in USEbench. We consider this a natural evolution or progression of the flurry of research proposing manifold different agents for different software engineering tasks. Existing specialized agents usually employ a fixed workflow and approach the given task in a few pre-defined steps. For example, AutoCodeRover tackles program repair tasks using a fixed two-phase workflow consisting of fault localization and patch generation. Now, what does it take to generalize an existing program repair agent, with fixed actions and pre-defined workflow, so that it can act as a unified software engineering agent (USEagent)? For each task or problem in USEbench, the agentic systems are only provided with a description (i.e., a bug report or the documentation of a method to generate) as well as access to a containerized environment of the project. As such, our meta-benchmark reveals inherent challenges regarding task-identification, workflow-configuration and measuring progress (esp. determining end-criteria).
@@
To address these requirements, we equip AutoCodeRover with a Meta-Agent, which is instructed to orchestrate the appropriate agents and construct a workflow on the fly. Over the course of this workflow, the Meta-Agent utilizes available actions to construct and maintain a project-state, constituting a structured consensus memory over the trajectories LLM components. Since OpenHands CodeActAgent is a general agent designed for solving general tasks, we consider it for baseline comparison, with no architectural changes applied.
@@ 2. Unified Software Engineering Benchmark
Table 1. Summary of primitive and compound tasks unified in USEbench
@@
To make progress towards building a unified agent for software engineering (USEagent), we first build a dataset of automated software engineering tasks. There is a rich environment of benchmarks, yet they are somewhat fragmented - they focus on specific types of software engineering tasks. Popular benchmarks encompass natural language issues for software (SWE-bench (Jimenez et al., 2024)), automated program repair (Defects4J (Just et al., 2014)), code generation (REPOCOD (Liang et al., 2024)), test generation (SWT (Mündler et al., 2024)) or documentation generation (CodeNet (Puri et al., 2021)), each accompanied by a metric to identify correct or optimal solutions. Next to their task, benchmarks often vary in scope too: the programming benchmarks often considered in foundational model evaluation sets (such as HumanEval (Chen et al., 2021b) , CodeNet (Puri et al., 2021), BigCode (Anonymous, 2021), EvalPlus (Liu et al., 2024)) cover method-snippets or class-level code that is evaluated against known examples. Benchmarks like Defects4J (Just et al., 2014) pose challenges for tools that require (correct) changes to a larger project, and are conceptually closer to experiences of developers. The recently proposed SWE-bench (Jimenez et al., 2024) offers a new angle on repository-scale changes by introducing natural language issues for program repair instead of a known test-suite. However, due to the natural language, some issues were too ambigious or un-solvable — spurring efforts of OpenAI to create a human-verified subset, called SWE-bench-verified (OpenAI, 2024a). Solving such natural language issues automatically brings the vision of a future AI software engineer closer to the capabilities of a human software engineer.
@@
In this paper, we propose USEbench, a unified software engineering benchmark as a meta-benchmark which moves beyond project-level code editing tasks suggested by SWE-bench. Our goal is to unify existing individual benchmark sets and design a comprehensive unified dataset capturing multiple SE tasks. Our proposal combines a set of existing benchmark sets, namely SWE-bench-verified (OpenAI, 2024a), SWT (Mündler et al., 2024), REPOCOD (Liang et al., 2024) and REPOTEST. Out of these SWE-bench-verified captures tasks which involve code editing or program modifications to achieve bug fixing / feature addition. SWT bench presents validation tasks for testing real-world code fixes in software projects. REPOCOD is a recent benchmark which captures code generation tasks in a software project. REPOTEST is a not previously published derivative of REPOCOD: While REPOCOD encompasses repository level code-generation after removing a method body; REPOTEST removes tests for a specified method-body and evaluation is based on code-coverage.
@@
The exact composition of USEbench from the constituent benchmarks shown in Table 1. Six different task types are covered, and for each task type several instances are included in USEbench. As a notable technical contribution, we provide a unified interface for agentic interaction with the unified benchmark. Previous benchmarks like SWE-bench provide a suitable interface for evaluation of code editing tasks i.e., after providing a .json file tasks are streamlined; still there is significant work involved in interacting with a software project, or extracting data from the constituent files of a software project. Thus, LLM agents proposed to work on SWE-bench tasks need to make these adjustments to work with SWE-bench, e.g. see (Orwall and contributors, 2025). We would want our unified benchmark USEbench to possess a greater degree of usability, so that LLM agents can be evaluated and compared in terms of their efficacy in conducting the software engineering tasks in the benchmark. To alleviate the burden on researchers, the constituent benchmarks are conveniently plugged behind a common interface revolving around docker images, providing options to read files and execute commands, regardless of the task. Due to the unified nature of our benchmark USEbench, we can also imagine unified software engineering agent (USEagent) working on the tasks in USEbench. Such a unified agent will be able to achieve complex software engineering tasks, which are solved via a sequence of the software engineering task types covered in the benchmark USEbench.
@@
We now show two concrete scenarios of more complex software engineering tasks constructed from primitive task types in USEbench, namely incomplete fix and feature development, elaborated in Table 1. To concretely see how the different task types in USEbench are combined into compound tasks, let us consider the scenario of handling an incomplete code fix. In such a scenario - we (i) initially test the code, find failing tests, (ii) then generate a fix for handling the failed tests (iii) test the fixed code and find more failed tests (iv) finally generate a fixed code which passes all tests. These four steps could be four different runs of a unified software engineering agent USEagent, and could combine different task types from USEbench, namely code fixing and test generation. The unified API of tasks in USEbench allows us to combine and alter tasks to capture more complex software engineering challenges, like incomplete fixes.
Another scenario of a complex software engineering task, and focus of both industry and research (Wang et al., 2024b), is the complete addition of a feature to a code-base. This includes a problem description, and requires a functional method-body as well as accommodating tests. How to solve the task, as well as any form of decomposition, re-iteration or co-evolution is left up to the users of the benchmark. The final evaluation of this derived scenario will be against a hidden test-suite, as well as the generated tests.
The GitHub repository of USEbench will be made publicly available soon.
@@ 3. Background: Agents for Software Engineering
Agents typically utilize LLMs for decision making and content generation, along with autonomously invoked tools to interact with external entities. These tools allow the agent access to knowledge beyond the LLMs’ training data, and allow the agent to influence the external environment through the tool interface. Agents often employ reasoning frameworks to orchestrate the LLM and the tool usage. One of the popular reasoning frameworks used in agents is ReAct (Yao et al., 2023), in which the LLMs are instructed to reason about output-traces and invoke tools in an interleaved manner.
@@
In software engineering, the target system that an agent interacts with is usually a software project. Another component of the external system is an execution environment, where software can be executed and tested. Existing software engineering agents (SE agents) proposed different ways to interact with the software project and the execution environment. AutoCodeRover (Zhang et al., 2024b) uses a set of project structure-aware tools (e.g. search_method_in_class to navigate the software codebase and gather necessary code context. It further performs test execution and spectrum-based localization (SBFL) (Wong et al., 2016) to pinpoint more relevant code locations. SWE-Agent (Yang et al., 2024) employs file-based tools to navigate files in the project, for example find_file, open, scroll_down, and together they form an agent-computer interface (Yang et al., 2024). OpenHands CodeActAgent (Wang et al., 2024b, a) provides a set of basic tools such as CmdRunAction to execute arbitrary bash commands inside a sandbox environment. A similar approach was followed by Google Jules agent (Google, 2024). Roughly summarized, the existing agents differ (intentionally) in how much structure and domain knowledge they introduce.
@@
Another aspect of agent design is the decision framework. Existing SE agents have explored options for more controlled decision-making beyond the ReAct framework, in the context of software engineering. AutoCodeRover (Zhang et al., 2024b) and Agentless (Xia et al., 2024) targets software maintenance tasks, and they break the agent’s execution into phases like context retrieval/localization and repair. Another example is RepairAgent (Bouzenia et al., 2024), which restricts available tools through a finite-state machine. The different designs in existing SE agents raise an important question – how much autonomy should be given to the agent during decision making? On one hand, agents can follow a fine-tuned workflow for the targeted software engineering (SE) task (e.g. in AutoCodeRover, Agentless, RepairAgent). In this design, agents operate in fixed phases (e.g. localization is the first phase for program repair), which makes them more cost-effective (Zhang et al., 2024b; Xia et al., 2024). With the task broken down into phases, it also becomes easier to leverage program analysis techniques in each of the phases. However, agents with this design have less autonomy, and as a result, they could not be directly applied to another kind of SE task. Applying a program repair agent to test generation would require drastic changes to its workflow. On the other hand, agents can be given more autonomy by having access to a general set of tools and the freedom to decide when to use these tools (e.g. in SWE-Agent and OpenHands CodeActAgent). This general design makes the agents applicable to many kinds of software engineering tasks. However, this generality makes it harder to exploit domain knowledge and program analysis in the agents. Moreover, when the task becomes complex and the agent execution gets longer, it is challenging to interpret the agent trajectory consisting of a list of general tool usages. In comparison, agents operating in “phases” are more interpretable, since we can inspect the results of each phase (e.g. what is the identified faulty location at the end of a localization phase).
@@ 4. Unified Software Engineering Agent
The first challenge is to adapt to tasks beyond a fixed workflow. Previous agents extract actions such as context retrieval or patch generation as clear “units of work” from developer workflows and organize them into state-machines or fixed workflows. This rigid organization of actions restricts their adaptability to other tasks. To achieve high agent autonomy we allow the actions to be freely composed and introduce a Meta-Agent that orchestrates the actions using a ReAct-style loop, forming a meta-layer over the actions, shown in Figure 1.
Figure 1. Concept: Meta-Agent abstracting over actions
@@
Starting from natural-language descriptions, the Meta-Agent invokes actions, observes the changes made by the action, reasons about the output, and decides on the next action to take. This framework enables on-the-fly composition of workflows, adjusting to different tasks, errors and changing states. ReAct-style decision making has been widely used in previous agents to orchestrate tools that directly interact with the codebase; which we extend towards action orchestration. Subsequent to the autonomy is the question of termination: A USEagent must be able to terminate ad libitum when the task is considered completed. We suggest a termination action, that the Meta-Agent can choose to invoke. The overall trajectory and choices of the Meta-Agent will construct an action-graph. In our action-graph, any node can be the start or end-point - for example, depending on the task, an action of “executing tests” can be the starting point or the last assurance before termination. The agent execution can be seen as an implicit construction of this action-graph on the fly.
@@
The next challenge is to provide the Meta-Agent with a set of reliable, yet flexible actions that are suitable for composition. Existing general agents such as OpenHands CodeActAgent (Wang et al., 2024b) and Google Jules (Google, 2024) provide the LLM with a console to execute any command. The action of executing a console command is extremely flexible, but hard to control. The agent’s execution consists of a sequence of console commands, which are difficult for human developers to interpret and raise (security) concerns. Thus, we propose to design our actions at a coarser granularity, where each action encapsulates a “unit of work” that developers carry out during the software engineering process. Each action receives instructions from the Meta-Agent on what outcomes it is supposed to achieve, rather than how it achieves them. The Meta-Agent and individual actions communicate through this intent-based interface. An example action is EditCode, where the Meta-Agent specifies what should be edited in the code, but not the exact details of making the edits. This design results in a set of modularized actions with clear responsibilities, that can be tested, extended and maintained individually. Moreover, when one existing action is improved, the improvements propagate across all involved task types. We present a sample list of actions for unified software engineering agents in Section 4.1.
@@
The third challenge is knowledge management: For complex tasks, the information provided a priori is not sufficient, and a central effort in solving tasks is identifying and producing relevant information. Once produced, this context should be made available to the agentic system, i.e. Meta-Agent and other actions. Existing work has defined this problem as memory management of agents (Zhang et al., 2024a; Han et al., 2024), where memory refers to historical information relevant to the current task. In literature on LLM agents, memory is often categorized into short-term (e.g., information within an ongoing conversation), long-term (e.g., historical data from previous sessions), and consensus memory (e.g., shared knowledge between multiple agents) (Han et al., 2024). In this work, we explicitly identify what information to memorize when designing a unified software engineering agent, and classify them into the different types of memory. We categorize the results of actions as short-term memory, which is only presented to the Meta-Agent to decide the next action. Project-specific knowledge, such as developer-written documentation, is considered long-term memory. These documents are embedded and stored in a vector database accessible for Retrieval-Augmented Generation (RAG) by the agent. Finally, artifacts generated by actions that are useful in guiding other actions constitute consensus memory. Such artifacts include program locations or code edits that can be notable to other actions. We introduce a structured task state that represents the consensus memory, which can be read and modified by different actions.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3773202isbn: 979-8-4007-2025-3/2026/04ccs: Software and its engineering Automatic programmingccs: Software and its engineering Software maintenance toolsccs: Software and its engineering Software evolution
@@ 1. Introduction
Large language models (LLMs) have shown promise in coding, reasoning, and problem solving. The model capability itself has improved significantly in terms of reasoning tasks, by training the models to always use reasoning paradigms such as chain of thoughts, without needing further prompting, as shown by GPT o1 (openai2024o1) and its descendant models such as o3. On top of improved foundation models, the software engineering community is distilling known best practices for downstream tasks into amplifiers for LLMs through agentic systems (yao2023react). An LLM agent for software engineering invokes various interfaces including testing and analysis tools autonomously driven by an LLM reasoning agent.
@@
While agentic systems inspire researchers (yao2023react; antoniades2024swesearch), start-ups (moatless_tools; AIDER2024) and major companies (wadhwa2024masai; rondon2025evaluating), we currently observe a strong specialization in the emerging technologies among the LLM agents for software. Agentless (xia2024agentless) performs program repair, Large Language Monkeys (brown2024large) produce unit tests, ExecutionAgent (bouzenia2024you) performs a project setup, and so forth. These specialized agents are reasonable efforts, especially as they require specialized metrics and datasets. However, with many agents comes the need to manage and maintain them as part of the future development environments!
@@
We propose a unified Software Engineering agent (USEagent) representing a consolidated agentic capability for software engineering tasks. Individual SE tasks could be coordinated to result in a more effective ensemble of tasks: better test-generation should benefit reproduction in program repair, review of patches should lead to better generated code, etc. With an integrated software engineering agent, it becomes much more feasible to co-opt it in future development environments (opinion25).
@@
As a foundation, we need a unified dataset of challenging SE tasks to test our unified agentic capabilities, for which we build Unified Software Engineering Bench, or USEbench for short. Recently, the SWE-bench dataset (jimenez2024swebench) has been proposed that captures a set of GitHub issues depicted in natural language and requires bug fixes or feature additions in software projects. Thus the SWE-bench dataset is replete with challenges in software maintenance for various real-life software projects. Our proposed benchmark USEbench is a meta-benchmark composing a diverse set of software engineering tasks (such as code generation, program repair, test generation) behind a unified application programming interface (API). Building a unified API is key, since it allows us to start thinking of a unified agentic capability which can handle different software engineering tasks. After constructing USEbench, we expand two popular agentic systems, one from industry and one from academia, AutoCodeRover (zhang2024autocoderover) and OpenHands CodeActAgent (wang2024openhands; wang2024executable) to solve the tasks in USEbench. We consider this a natural evolution or progression of the flurry of research proposing manifold different agents for different software engineering tasks. Existing specialized agents usually employ a fixed workflow and approach the given task in a few pre-defined steps. For example, AutoCodeRover tackles program repair tasks using a fixed two-phase workflow consisting of fault localization and patch generation. Now, what does it take to generalize an existing program repair agent, with fixed actions and pre-defined workflow, so that it can act as a unified software engineering agent (USEagent)? For each task or problem in USEbench, the agentic systems are only provided with a description (i.e., a bug report or the documentation of a method to generate) as well as access to a containerized environment of the project. As such, our meta-benchmark reveals inherent challenges regarding task-identification, workflow-configuration and measuring progress (esp. determining end-criteria).
@@
To address these requirements, we equip AutoCodeRover with a Meta-Agent, which is instructed to orchestrate the appropriate agents and construct a workflow on the fly. Over the course of this workflow, the Meta-Agent utilizes available actions to construct and maintain a project-state, constituting a structured consensus memory over the trajectories LLM components. Since OpenHands CodeActAgent is a general-purpose agent designed to solve a variety of tasks, we consider it for baseline comparison, with no architectural changes applied.
@@ 2. Unified Software Engineering Benchmark
Table 1. Summary of primitive and compound tasks unified in USEbench.
@@
To make progress towards building a unified agent for software engineering (USEagent), we first build a dataset of automated software engineering tasks. There is a rich environment of benchmarks, yet they are somewhat fragmented - they focus on specific types of software engineering tasks. Popular benchmarks encompass natural language issues for software (SWE-bench (jimenez2024swebench)), automated program repair (Defects4J (just2014defects4j)), code generation (REPOCOD (liang2024can)), test generation (SWT (mundler2024swtbench)) or documentation generation (CodeNet (CodeNet)), each accompanied by a metric to identify correct or optimal solutions. Next to their task, benchmarks often vary in scope too: the programming benchmarks (such as HumanEval (HumanEval) , CodeNet (CodeNet), BigCode (BigCodeProject), EvalPlus (liu2024your)) cover method-snippets or class-level code that is evaluated against known examples. The recently proposed SWE-bench (jimenez2024swebench) offers a new angle on repository-scale changes by introducing natural language issues for program repair instead of a known test-suite. However, due to the natural language, some issues were too ambigious or un-solvable — spurring efforts of OpenAI to create a human-verified subset, called SWE-bench-verified (OpenAISWEBench2023). Solving such natural language issues automatically brings the vision of a future AI software engineer closer to the capabilities of a human software engineer.
@@
In this paper, we propose USEbench, a unified software engineering benchmark as a meta-benchmark which moves beyond project-level code editing tasks suggested by SWE-bench. Our goal is to unify existing individual benchmark sets and design a comprehensive unified dataset capturing multiple SE tasks. Our proposal combines a set of existing benchmark sets, namely SWE-bench-verified (OpenAISWEBench2023), SWT (mundler2024swtbench), REPOCOD (liang2024can) and REPOTEST. Out of these SWE-bench-verified captures tasks which involve code editing or program modifications to achieve bug fixing / feature addition. SWT bench presents validation tasks for testing real-world code fixes in software projects. REPOCOD is a recent benchmark which captures code generation tasks in a software project. REPOTEST is a not previously published derivative of REPOCOD: While REPOCOD encompasses repository level code-generation after removing a method body; REPOTEST removes tests for a specified method-body and evaluation is based on code-coverage.
@@
The exact composition of USEbench from the constituent benchmarks is shown in Table 1. Six different task types are covered, and for each task type several instances are included in USEbench. As a notable technical contribution, we provide a unified interface for agentic interaction with the unified benchmark. Previous benchmarks like SWE-bench provide a suitable interface for evaluation of code editing tasks i.e., after providing a .json file tasks are streamlined; still there is significant work involved in interacting with a software project, or extracting data from the constituent files of a software project. Thus, LLM agents proposed to work on SWE-bench tasks need to make these adjustments to work with SWE-bench, e.g. see (moatless_tools). We would want our unified benchmark USEbench to possess a greater degree of usability, so that LLM agents can be evaluated and compared in terms of their efficacy in conducting the software engineering tasks in the benchmark. To alleviate the burden on researchers, the constituent benchmarks are conveniently plugged behind a common interface revolving around docker images, providing options to read files and execute commands, regardless of the task. Due to the unified nature of our benchmark USEbench, we can also imagine unified software engineering agent (USEagent) working on the tasks in USEbench. Such a unified agent will be able to achieve complex software engineering tasks, which are solved via a sequence of the software engineering task types covered in the benchmark USEbench.
@@
We now show two concrete scenarios of more complex software engineering tasks constructed from primitive task types in USEbench, namely incomplete fix and feature development, elaborated in Table 1. To concretely see how the different task types in USEbench are combined into compound tasks, let us consider the scenario of handling an incomplete code fix. In such a scenario, we (i) test the code and find failing tests, (ii) generate a fix for the failed tests, (iii) test the fixed code and find more failed tests, and (iv) finally generate a fixed code which passes all tests. These four steps could be four different runs of a software engineering agent, and could combine different task types from USEbench, namely code fixing and test generation. The unified API of tasks in USEbench allows us to combine and alter tasks to capture more complex software engineering challenges, like incomplete fixes. Another scenario of a complex software engineering task is the complete addition of a feature to a codebase. This includes a problem description, and requires a functional method-body as well as accommodating tests. How to solve the task, as well as any form of decomposition, re-iteration or co-evolution is left up to the users of the benchmark. The final evaluation of this derived scenario will be against a hidden test-suite, as well as the generated tests.
@@ 3. Background: Agents for Software Engineering
Agents typically utilize LLMs for decision making and content generation, along with autonomously invoked tools to interact with external entities. These tools allow the agent access to knowledge beyond the LLMs’ training data, and allow the agent to influence the external environment through the tool interface. Agents often employ reasoning frameworks to orchestrate the LLM and the tool usage. One of the popular reasoning frameworks used in agents is ReAct (yao2023react), in which the LLMs are instructed to reason about output-traces and invoke tools in an interleaved manner.
@@
In software engineering, the target system that an agent interacts with is usually a software project. Another component of the external system is an execution environment, where software can be executed and tested. Existing software engineering agents (SE agents) proposed different ways to interact with the software project and the execution environment. AutoCodeRover (zhang2024autocoderover) uses a set of project structure-aware tools (e.g. search_method_in_class to navigate the software codebase and gather necessary code context. It further performs test execution and spectrum-based localization (SBFL) (wong2016survey) to pinpoint more relevant code locations. SWE-Agent (yang2024swe) employs file-based tools to navigate files in the project, for example find_file, open, scroll_down, and together they form an agent-computer interface (yang2024swe). OpenHands CodeActAgent (wang2024openhands; wang2024executable) provides a set of basic tools such as CmdRunAction to execute arbitrary bash commands inside a sandbox environment. A similar approach was followed by Google Jules agent (google2024gemini2). Roughly summarized, the existing agents differ (intentionally) in how much structure and domain knowledge they introduce.
@@
Another aspect of agent design is the decision framework. Existing SE agents have explored options for more controlled decision-making beyond the ReAct framework, in the context of software engineering. AutoCodeRover (zhang2024autocoderover) and Agentless (xia2024agentless) targets software maintenance tasks, and they break the agent’s execution into phases like context retrieval/localization and repair. Another example is RepairAgent (bouzenia2024repairagent), which restricts available tools through a finite-state machine. The different designs in existing SE agents raise an important question – how much autonomy should be given to the agent during decision making? On one hand, agents can follow a fine-tuned workflow for the targeted software engineering (SE) task (e.g. in AutoCodeRover, Agentless, RepairAgent). In this design, agents operate in fixed phases (e.g. localization is the first phase for program repair), which makes them more cost-effective (zhang2024autocoderover; xia2024agentless). With the task broken down into phases, it also becomes easier to leverage program analysis techniques in each of the phases. However, agents with this design have less autonomy, and as a result, they could not be directly applied to another kind of SE task. Applying a program repair agent to test generation would require drastic changes to its workflow. On the other hand, agents can be given more autonomy by having access to a general set of tools and the freedom to decide when to use these tools (e.g. in SWE-Agent and OpenHands CodeActAgent). This general design makes the agents applicable to many kinds of software engineering tasks. However, this generality makes it harder to exploit domain knowledge and program analysis in the agents. Moreover, when the task becomes complex and the agent execution gets longer, it is challenging to interpret the agent trajectory consisting of a list of general tool usages. In comparison, agents operating in “phases” are more interpretable, since we can inspect the results of each phase (e.g. what is the identified faulty location at the end of a localization phase).
@@ 4. Unified Software Engineering Agent
The first challenge is to adapt to tasks beyond a fixed workflow. Previous agents extract actions such as context retrieval or patch generation as clear “units of work” from developer workflows and organize them into state-machines or fixed workflows. This rigid organization of actions restricts their adaptability to other tasks. To achieve high agent autonomy we allow the actions to be freely composed and introduce a Meta-Agent that orchestrates the actions using a ReAct-style loop, forming a meta-layer over the actions, shown in Figure 1. Starting from natural-language descriptions, the Meta-Agent invokes actions, observes the changes made by the action, reasons about the output, and decides on the next action to take. This framework enables on-the-fly composition of workflows, adjusting to different tasks, errors and changing states. ReAct-style decision making has been widely used in previous agents to orchestrate tools that directly interact with the codebase; which we extend towards action orchestration. The overall trajectory and choices of the Meta-Agent will construct an action-graph. In our action-graph, any node can be the start or end-point - for example, depending on the task, an action of “executing tests” can be the starting point or the last assurance before termination. The agent execution can be seen as an implicit construction of this action-graph on the fly.
Figure 1. Concept: Meta-Agent abstracting over actions.
@@
The next challenge is to provide the Meta-Agent with a set of reliable, yet flexible actions that are suitable for composition. Existing general agents such as OpenHands CodeActAgent (wang2024openhands) and Google Jules (google2024gemini2) provide the LLM with a console to execute any command. The action of executing a console command is extremely flexible, but hard to control. The agent’s execution consists of a sequence of console commands, which are difficult for human developers to interpret and raise security concerns. Thus, we propose to design our actions at a coarser granularity, where each action encapsulates a “unit of work” that developers carry out during the software engineering process. Each action receives instructions from the Meta-Agent on what outcomes it is supposed to achieve, rather than how it achieves them. The Meta-Agent and individual actions communicate through this intent-based interface. An example action is EditCode, where the Meta-Agent specifies what should be edited in the code, but not the exact details of making the edits. This design results in a set of modularized actions with clear responsibilities, that can be tested, extended and maintained individually. Moreover, when one action is improved, the improvements propagate across all involved task types. We present a sample list of actions for unified software engineering agents in Section 4.1.
@@
The third challenge is knowledge management: For complex tasks, the information provided a priori is not sufficient, and a central effort in solving tasks is identifying and producing relevant information. Once produced, this context should be made available to the agentic system, i.e. Meta-Agent and other actions. Existing work has defined this problem as memory management of agents (zhang2024survey; han2024llm), where memory refers to historical information relevant to the current task. In literature on LLM agents, memory is often categorized into short-term (e.g., information within an ongoing conversation), long-term (e.g., historical data from previous sessions), and consensus memory (e.g., shared knowledge between multiple agents) (han2024llm). In this work, we explicitly identify what information to memorize when designing a unified software engineering agent, and classify them into the different types of memory. We categorize the results of actions as short-term memory, which is only presented to the Meta-Agent to decide the next action. Project-specific knowledge, such as developer-written documentation, is considered long-term memory. These documents are embedded and stored in a vector database accessible for Retrieval-Augmented Generation (RAG) by the agent. Finally, artifacts generated by actions that are useful in guiding other actions constitute consensus memory. Such artifacts include program locations or code edits that can be notable to other actions. We introduce a structured task state that represents the consensus memory, which can be read and modified by different actions.